OpenStudy (anonymous):

A women bought some large flowers for 18 dollars each and some frames for 5 dollars each at a closeout sale. If she bought 23 frames for 206 dollars, find how many of each type she bought

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

I am having trouble understanding this question. Did she buy some frames and flowers and spend $206 ? Because 23 frames, at $5 each, it would cost $115. I am going to assume you meant she bought both for 206.If that is the case, she spent 115 dollars on frames. Leaving (206 - 115) $91. If flowers cost $15, then she would have to buy (91/15) = 6.06 rounds to 6. So, she bought 23 frames at 115, and 6 flowers at 90 totaling $205. She would get $1 back. Unless of course there is tax, then she would owe more.

OpenStudy (anonymous):

@ParthKohli .....did I do something wrong or is some info missing from this problem.....or could it just be worded wrong ?

Parth (parthkohli):

OK, you assume that the number of frames you got is \(x\). Then you got \(15 -x\) flowers. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It says she bought 23 frames for 206, but frames are $5 each and that doesn't equal. Did she buy flowers and frames for 206 ?

OpenStudy (anonymous):

if that is the case, an equation can be set up...right ? oops...I read the problem wrong....flowers are $18. 18f + 115 = 206 would this be a correct equation

OpenStudy (anonymous):

did she buy 23 frames or 23 items ?

OpenStudy (anonymous):

@ParthKohli ...are you still there ? I don't know why I am worried about this...the one that posted is not even online.

Parth (parthkohli):

Whoops, I was away.

Parth (parthkohli):

23 items, yes.

OpenStudy (anonymous):

that changes everything

Parth (parthkohli):

We have the total cost given as,\[5(x) + 18(23 - x) = 206\]

OpenStudy (anonymous):

got it......thanks again.....you should be a teacher :)

Parth (parthkohli):

hehe